Heathcote Corporation is a manufacturer that uses job-order costing. The company closes out any overapplied or underapplied overhead to Cost of Goods Sold at the end of the year. The company has supplied the following data for the just completed year:

Estimated total manufacturing overhead at the beginning of the year $ 546,000
Estimated direct labor-hours at the beginning of the year 42,000 direct labor-hours

Results of operations:

Actual direct labor-hours 47,000 direct labor-hours
Manufacturing overhead:
Indirect labor cost $ 152,000
Other manufacturing overhead costs incurred $ 454,000
Cost of goods manufactured $ 1,569,000
Cost of goods sold (unadjusted) $ 1,458,000

The total amount of manufacturing overhead applied to production is:
Answer #1

Predetermined Overhead rate = Estimated overhead/Estimated total labor hours

= 546,000/42,000 = 13 per direct labor hour

Overhead applied = Actual direct labor hours * 13

= 47,000 * 13

= 611,000